40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 13 22:32:54 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered PMC_LO.pngIMG_2093.JPGIMG_2091.JPGIMG_2095.JPGIMG_2096.JPG
    Reply  Sun Nov 15 22:43:48 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered C1-MULTI_E6875C_TIMESERIES-1131408017-86400.pngC1-MULTI_E6875C_TIMESERIES-1131580817-86400.png
       Reply  Mon Nov 16 19:05:59 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
          Reply  Mon Nov 16 21:32:49 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
             Reply  Tue Nov 17 16:21:10 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
                Reply  Wed Nov 18 16:51:58 2015, Koji, Summary, PSL, PMC servo calibration PMC_err_cal.pdf
                   Reply  Wed Nov 18 16:53:17 2015, Koji, Summary, PSL, PMC PZT capacitance 
                      Reply  Wed Nov 18 17:09:22 2015, Koji, Summary, PSL, PMC Servo analysis PMC_OLTF.pdfPMC_DSP.pdf
                         Reply  Fri Nov 20 15:44:12 2015, Koji, Summary, PSL, Added 17.5kHz LPF to the PMC servo PMC_OLTF.pdfPMC_noise_comparison.pdf
Message ID: 11768     Entry time: Mon Nov 16 19:05:59 2015     In reply to: 11765     Reply to this: 11769
Author: Koji 
Type: Summary 
Category: PSL 
Subject: PMC servo circuit review, follow up measurements 

PMC follow up measurements have been done. The servo circuit was reviewed.

Now the PMC, IMC, X/Y arms are locked and aligned waiting for the IFO work although I still think something is moving (ITMX?)
as the FPMI fringe is quite fast.

ELOG V3.1.3-